OpenStudy (anonymous):

Solve for b. d = 3a + 3b A. b = d-3a/3 B. b = 3a-d/3 C. b = d-3a/a D. b = d-3a

OpenStudy (anonymous):

jst solve for b 3b=d-3a b=(d-3a)/3

OpenStudy (anonymous):

Firstly, we must put all foreign terms that arent B to the other side of the equation, to do this we have to remove the 3a. To move the 3a to the other side we simply subtract it, once its on the other side it becomes 3a - d= 3b

OpenStudy (anonymous):

@Tcookie009 : get clear with the options as there should be diffrnc between a and c